OpenStudy (anonymous):

6 3/8 divided by 2 3/7??

OpenStudy (anonymous):

6 3/8 = (6*8+3)/8 = 51/8 2 3/7 = (2*7+3)/7 = 17/7 \[\frac{\frac{51}{8}}{\frac{17}{7}}=\frac{51}{8}*\frac{7}{17}=\frac{17*3}{8}*\frac{7}{17}=\frac{3}{8}*7=\frac{21}{8}\]

OpenStudy (anonymous):

21/8 = 16 5/8